A balanced Y load having a 15+j25 Ω load impedance in each leg is connected to a three-phase, four-wire, Y-connected generator having a line voltage of 230 V. The line impedance per phase is 0.5+ j1.5 Ω . Answer the following:

a. Draw the circuit showing the given values.

b. Calculate the line voltages of the load.

c. Calculate the phase currents of the load.

d. Calculate the power consumed by the 3-phase load.



1
Expert's answer
2021-01-06T06:04:42-0500

The image attached shows the schematic.




b) Phase voltage of source

"=\frac{230}{\sqrt{3}} =132.79 V."

Phase voltage on load side"=132.79 \times \frac{15+j25}{(15+j25) +(0.5+j1.5) }=126.1-j1.41=126.1\angle-0.64\degree"

Hence line voltage on load"=126.1\times \sqrt{3}= 218.41V"


C) phase current of the load "=\frac{132.79}{(15+j25) +(0.5+j1.5) } = 4.32\angle-59.68A"


d) power consumed by 3 phase load="3\times i^2\times Rload=3\times4.32^2\times15=839.81W"